# Self-Consistency Preference Optimization

### v2

@@ 1 Introduction
To address this issue, we introduce Self-consistency Preference Optimization (ScPO). ScPO is an approach to self-train LLMs for complex problem-solving tasks without access to gold solutions or final answers in the training data. Our approach leverages the concept of self-consistency (Wang et al., 2023), an inference-time only approach that improves performance on reasoning tasks by generating multiple solutions using the LLM and choosing the most frequent final answer. More consistent answers are more likely to be correct because mistakes made by the model are often random, so incorrect solutions are unlikely to lead to the same answer multiple times (Fischler & Bolles, 1981; Chen et al., 2023). In ScPO, the self-consistency concept is instead applied during unsupervised self-training. The method consists of (i) selecting model-generated queries, (ii) annotating preference pairs using the most self-consistent response (winner) and least self-consistent response (loser), and (iii) optimizing a loss function that is weighted for each instance depending on the model’s confidence in the preference pair. Additionally, we propose a semi-supervised variant of ScPO that jointly trains LLMs on labeled and unlabeled instances, taking advantage of human annotations whenever available. Unlike self-consistency applied at inference time, ScPO does not increase inference-time compute, but they can also be combined together for better performance.
@@ 2 Self-consistency Preference Optimization
As depicted in Figure 1, ScPO is an unsupervised iterative training method that starts with a base language model. Each iteration makes use of existing training problems/queries (without labels) as well as newly generated problems. The self-consistency metric is used in both generating new problems and building preference pairs. We describe each step of ScPO’s iterative training setup below. All prompts for solution generation and new problem generation can be found in Appendix C.
@@
ScPO assumes access to an initial base model $M_{0}$ and a small amount of (seed) high-quality unlabeled queries, which are typically complex reasoning problems. The model will be trained and updated at each training iteration resulting in models $M_{1},M_{2},\cdots,M_{T}$, where $T$ is the total number of iterations. Instead of gold labels (answers) for responses, ScPO uses the consistency of the model $M_{t}$, as measured by a real-valued vote function $\mathcal{V}(\cdot)$ defined below, to rate and rank the quality of each response. Our vote function is based on self-consistency (Wang et al., 2023) of the model ScPO can also be used with any measure of model consistency such as internal consistency (Liang et al., 2024) or universal consistency (Chen et al., 2024a).
@@
For each problem $x$ in the training data $\mathcal{D}_{t}$, we use temperature-based sampling with the current model $M_{t}$ to generate $k$ responses $\bar{\bm{y}}_{x}=\{y_{1},y_{2},\cdots,y_{k}\}$ sampled from $M_{t}(\cdot|x)$ including any rationales, e.g., chain-of-thought (Wei et al., 2022), followed by the final answer. Following Wang et al. (2023), the vote function $\mathcal{V}(\cdot)$ extracts the final answer corresponding to each response $y\in\bar{{\bm{y}}}_{x}$ via $\mathrm{ans}(\cdot)$ and returns the relative frequency of the final answer, i.e., $\mathcal{V}(y)\!=\!\sum_{m=1}^{k}\mathbbm{1}(\mathrm{ans}(y_{m})\!=\!\mathrm{ans}(y))$. As illustrated in Figure 1 (middle), using the vote function, we create preference pairs $\mathcal{D}_{t}^{\mathrm{pairs}}$ by selecting the most consistent response as the chosen (winning) response and selecting the least consistent one as the rejected (losing) response, provided that the vote of the chosen response is greater than a threshold $\tau$ [^1] By design, several responses can share a final answer (but for example, their chain-of-thought may be different). So, we cluster the responses by final answer and pick a response at random among them. In other words,
$\mathcal{D}_{t}^{\mathrm{pairs}}=\{(x,y^{+},y^{-})|x\in\mathcal{D}_{t},y^{+}\!=\!\arg\max_{y\in\bar{{\bm{y}}}_{x}}\mathcal{V}(y),\ y^{-}\!=\!\arg\min_{y\in\bar{{\bm{y}}}_{x}}\mathcal{V}(y),\text{ and }\mathcal{V}(y^{+})\geq\tau\}.$
@@
ScPO operates under the assumption that when multiple responses sampled for problem $x$ map to the same answer, then the predicted answer is likely to be correct, the same assumption as in Wang et al. (2023). Consequently, we use consistency via a vote function $\mathcal{V}(\cdot)$ as a proxy to create preference pairs. However, at the same time, the number of votes attained by a response can also reflect the model’s confidence in the response (Xiong et al., 2024; Kabra et al., 2024), implying that pairs where the vote margin – the difference in votes attained by the chosen vs. the rejected response – is larger, are of higher quality and vice-versa (refer to Appendix A). We model this in ScPO’s training by using an instance-level weight $w(x)$ to the loss, i.e., for the preference pair $(x,y^{+},y^{-})\in\mathcal{D}_{t}^{\mathrm{pairs}}$, $w(x)\!=\!\big{(}\mathcal{V}(y^{+})-\mathcal{V}(y^{-})\big{)}/k$, where $k$ is the total number of responses generated for each question (total number of votes cast). [^2] This normalization makes sure the weights are always between 0 and 1. We thus use the following loss function: $\mathcal{L}_{\textsc{ScPO}{}}(y^{+},y^{-}|x)\!$ $=\!\underbrace{-w(x)\log\sigma\left(\beta\log\frac{M_{\theta}(y^{+}\!\mid\!x)}{M_{t}(y^{+}\mid x)}-\beta\log\frac{M_{\theta}(y^{-}\!\mid\!x)}{M_{t}(y^{-}\!\mid x)}\right)}_{\text{Weighted DPO Loss}}\underbrace{-\frac{\alpha w(x)}{|y^{+}|}\log M_{\theta}(y^{+}\!\mid\!x)}_{\text{Weighted NLL Loss}}.$ The loss includes a DPO and NLL term similar to the recently introduced supervised IRPO (Pang et al., 2024) loss, but in our case we have an unsupervised objective and use our introduced weighted loss. Here $\sigma(\cdot)$ denotes the sigmoid function, and $\alpha,\beta$ are hyperparameters of the loss function, and $\theta$ represents the LLM parameters being trained in the current iteration. At the $t^{\text{th}}$ iteration, we use the initialized model $M_{t}$ as the reference model in the DPO loss (Rafailov et al., 2024). After training on this loss, the trained model is used to initialize the next iteration, i.e., $M_{t+1}\leftarrow M_{\theta}$.
@@ 3 Experimental Setup
We evaluate the effectiveness of ScPO on a range of mathematical and logical reasoning datasets:
@@
We compare models trained with ScPO in unsupervised (denoted as ScPOUnsup.) and semi-supervised (denoted as ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$) settings against the following baselines:
@@
llr@ / lcc Method Train Data (K) Test Acc. (%)
# Seed Gen. Greedy SC (8-way)
@@
 IRPORM iter $M_{1}$ 5.5 - 48.67 69.98
iter $M_{2}$ 4.4 - 50.11 61.25
 ScPOUnsup. iter $M_{1}$ 5.3 - 61.03 71.49
iter $M_{2}$ 1.4 5.1 63.91 71.11
@@
 IRPOGold iter $M_{1}$ 4.4 - 61.41 72.93
iter $M_{2}$ 5.7 - 64.29 72.56
 ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ iter $M_{1}$ 4.4 1.9 63.61 74.30
iter $M_{2}$ 5.7 4.5 66.64 74.75
Table 1: GSM8K zero-shot accuracy after training Llama-3 Base 8B with ScPO and baselines, using greedy or self-consistency (SC)-based inference. The best performance is in bold, and second-best is underlined. We list train set sizes for each method: “Seed” corresponds to seed problems in the train set, whereas “Gen.” indicates additional problems generated by the model (without answers). IRPOGold, and ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$, highlighted in green, use the gold answers to create preference pairs (when available, indicated with ).
@@ 4.1 Math Reasoning
Comparing methods on GSM8K, in Section 3, we observe that training with only one iteration of ScPO outperforms the zero-shot seed model and IRPORM, by $22.74\%$ and $12.36\%$, respectively, using greedy decoding. Similarly, on MATH (cf. Section 4.1), two iterations of ScPOUnsup. yields an improvement of $5.26\%$ and $1.64\%$ respectively compared to the same two baselines. We further note that while IRPORM is not given direct access to the gold labels, it uses the ArmoRM, which has been trained on human-annotated step-level data based on MATH’s train set (Lightman et al., 2023; Wang et al., 2024a). Hence, ScPO’s improvement over IRPORM would likely be larger if the RM had not used in-domain gold labels during training. Overall, we find ScPO has the ability to outperform RMs, especially in out-of-distribution settings.
@@
llr@ / lcc Method Train Data (K) Test Acc. (%)
# Seed Gen. Greedy SC (8-way)
@@
 IRPORM iter $M_{1}$ 6.4 - 18.06 24.20
iter $M_{2}$ 6.5 - 18.08 22.64
 ScPOUnsup. iter $M_{1}$ 0.6 1.2 17.36 25.70
iter $M_{2}$ 1.2 2.5 19.72 24.58
@@
 IRPOGold iter $M_{1}$ 2.7 - 18.64 26.88
iter $M_{2}$ 3.0 - 20.32 26.88
 ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ iter $M_{1}$ 2.7 1.2 19.88 27.35
iter $M_{2}$ 3.0 2.2 20.48 26.92
Table 2: MATH zero-shot accuracy after training Llama-3 Base 8B with ScPO and baselines, using greedy or self-consistency (SC)-based inference. Train data size: “Seed” corresponds to seed queries in the train set, “Gen.” are additional model-generated problems (without answers). IRPOGold and ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$, highlighted in green, use gold answers to train (indicated with ).
Iterations of ScPO improve reasoning.
Unsupervised ScPO is comparable to IRPO training with gold labels.
We can compare the unsupervised training of ScPO with the supervised training using gold labels of IRPO in Sections 3 and 4.1. The results show that ScPOUnsup. without using any gold labels can yield comparable accuracy to IRPOGold on GSM8K and MATH with $<1\%$ gap in greedy performance and $<2\%$ gap in accuracy using 8-way self-consistency after two iterations of training ($M_{2}$). This comparable performance of ScPOUnsup. is likely due to high correlation (0.8 across the datasets) between the vote shares and accuracy on the test set, as further discussed in Appendix A. Note that on tasks that are challenging for the seed model $M_{0}$, such as MATH, we can only bootstrap a small set of examples from the original set of training problem as compared to IRPO (i.e., only around a quarter of examples obtain a clear majority answer). However, we can offset this gap in training data by generating new problems using few-shot prompting (cf. Section 2) and creating preference pairs using our self-consistency method. This helps provide improvements during the second iteration.
Semi-supervised training with ScPO outperforms IRPO.
Lastly, in Sections 3 and 4.1, we evaluate the semi-supervised version of ScPO combined with using gold labels. We find that on GSM8K, ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ improves the greedy accuracy by $2.35\%$ and SC accuracy by $2.19\%$ in comparison to IRPOGold. Similar trends hold on the MATH dataset, where one iteration of ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ outperforms IRPOGold by $1.24\%$ using greedy decoding. These results show the utility of using ScPO to bootstrap from model-generated problems even with access to a labeled training set.
@@
$M_{1}$ w/ ScPOUnsup. 0.4 / 1.0 17.0 54.3 2.5 47.6
$M_{2}$ w/ ScPOUnsup. 0.4 / 2.2 18.1 58.2 2.5 45.2
Table 3: ZebraLogic test performance after unsupervised training of Llama-3 Instruct 8B with ScPO, compared to baselines. “Seed” corresponds to original puzzles in the test set, whereas “Gen.” indicates additional puzzles generated. Model’s performance reported from the Leaderboard.

### v3

@@ 1 Introduction
To address this, we introduce Self-consistency Preference Optimization (ScPO). ScPO is an approach to self-train LLMs for complex problem-solving tasks without access to gold solutions or final answers in the training data. Our approach leverages the concept of self-consistency (Wang et al., 2023), an inference-time only approach that improves performance on reasoning tasks by generating multiple solutions using the LLM and choosing the most frequent final answer. More consistent answers are more likely to be correct because mistakes made by the model are often random, so incorrect solutions are unlikely to lead to the same answer multiple times (Fischler & Bolles, 1981; Chen et al., 2023). In ScPO, the self-consistency concept is instead applied during unsupervised self-training. The method consists of (i) selecting model-generated queries, (ii) annotating preference pairs using the most self-consistent response (winner) and least self-consistent response (loser), and (iii) optimizing a loss function that is weighted for each instance depending on the model’s confidence in the preference pair. Additionally, we propose a semi-supervised variant of ScPO that jointly trains LLMs on labeled and unlabeled instances, taking advantage of human annotations whenever available. Unlike self-consistency applied during inference, ScPO does not increase inference-time compute, but they can also be combined together for better performance.
@@ 2 Self-consistency Preference Optimization
As depicted in Figure 1, ScPO is an unsupervised iterative training method that starts with a base language model. Each iteration makes use of existing training problems/queries (without labels) as well as newly generated problems. The self-consistency metric is used in both generating new problems and building preference pairs. We describe each step of ScPO’s iterative training setup below. All prompts for solution generation and new problem generation can be found in Appendix D.
@@
ScPO assumes access to an initial base model $M_{0}$ and a small amount of (seed) high-quality unlabeled queries, which are typically complex reasoning problems. The model will be trained and updated at each training iteration resulting in models $M_{1},M_{2},\cdots,M_{T}$, where $T$ is the total number of iterations. Instead of gold labels (answers) for responses, ScPO uses the consistency of the model $M_{t}$, as measured by a real-valued vote function $\mathcal{V}(\cdot)$ defined below, to rate and rank the quality of each response. Our vote function is based on self-consistency (Wang et al., 2023) of the model. In fact, ScPO can also be used with any measure of model consistency such as internal consistency (Liang et al., 2024) or universal consistency (Chen et al., 2024a).
@@
For each problem $x$ in the training data $\mathcal{D}_{t}$, we use temperature-based sampling with the current model $M_{t}$ to generate $k$ responses $\bar{\bm{y}}_{x}=\{y_{1},y_{2},\cdots,y_{k}\}$ sampled from $M_{t}(\cdot|x)$ including any rationales, e.g., chain-of-thought (Wei et al., 2022), followed by the final answer. Following Wang et al. (2023), the vote function $\mathcal{V}(\cdot)$ extracts the final answer corresponding to each response $y\in\bar{{\bm{y}}}_{x}$ via $\mathrm{ans}(\cdot)$ and returns the relative frequency of the final answer, i.e., $\mathcal{V}(y)\!=\!\sum_{m=1}^{k}\mathbbm{1}(\mathrm{ans}(y_{m})\!=\!\mathrm{ans}(y))$. As illustrated in Figure 1 (middle), using the vote function, we create preference pairs $\mathcal{D}_{t}^{\mathrm{pairs}}$ by selecting the most consistent response as the chosen (winning) response and selecting the least consistent one as the rejected (losing) response, provided that the vote of the chosen response is greater than a threshold $\tau$ [^1] By design, several responses can share a final answer (but for example, their chain-of-thought may be different). So, we cluster the responses by final answer and pick a response at random. In other words,
$\mathcal{D}_{t}^{\mathrm{pairs}}$ =$=\{(x,y^{+},y^{-})\mid x\in\mathcal{D}_{t},y^{+}\!=\!\arg\max_{y\in\bar{{\bm{y}}}_{x}}\mathcal{V}(y),$
$\qquad y^{-}\!=\!\arg\min_{y\in\bar{{\bm{y}}}_{x}}\mathcal{V}(y),\text{ and }\mathcal{V}(y^{+})\geq\tau\}.$
@@
ScPO operates under the assumption that when multiple responses sampled for problem $x$ map to the same answer, then the predicted answer is likely to be correct, the same assumption as in Wang et al. (2023). Consequently, we use consistency via a vote function $\mathcal{V}(\cdot)$ as a proxy to create preference pairs. However, at the same time, the number of votes attained by a response can also reflect the model’s confidence in the response (Xiong et al., 2024; Kabra et al., 2024), implying that pairs where the vote margin – the difference in votes attained by the chosen vs. the rejected response – is larger, are of higher quality and vice-versa (refer to Appendix A). We model this in ScPO’s training by using an instance-level weight $w(x)$ to the loss, i.e., for the preference pair $(x,y^{+},y^{-})\in\mathcal{D}_{t}^{\mathrm{pairs}}$, $w(x)\!=\!\big{(}\mathcal{V}(y^{+})-\mathcal{V}(y^{-})\big{)}/k$, where $k$ is the total number of responses generated for each question (total number of votes cast). [^2] This normalization ensures that weights $w(x)\in[0,1]$. We thus use the following loss function:
$\mathcal{L}_{\textsc{ScPO}{}}(y^{+},y^{-}|x)=$
$\ \underbrace{-w(x)\log\sigma\left(\beta\log\frac{M_{\theta}(y^{+}\!\mid\!x)}{M_{t}(y^{+}\mid x)}-\beta\log\frac{M_{\theta}(y^{-}\!\mid\!x)}{M_{t}(y^{-}\!\mid x)}\right)}_{\text{Weighted DPO Loss}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\ \underbrace{-~{}~{}~{}\frac{\alpha w(x)}{|y^{+}|}\log M_{\theta}(y^{+}\!\mid\!x)}_{\text{Weighted NLL Loss}}.$
The loss includes a DPO and NLL term similar to the recently introduced supervised IRPO (Pang et al., 2024) loss, but in our case we have an unsupervised objective and use our introduced weighted loss. Here $\sigma(\cdot)$ denotes the sigmoid function, and $\alpha,\beta$ are hyperparameters of the loss function, and $\theta$ represents the LLM parameters being trained in the current iteration. At the $t^{\text{th}}$ iteration, we use the initialized model $M_{t}$ as the reference model in the DPO loss (Rafailov et al., 2024). After training on this loss, the trained model is used to initialize the next iteration, i.e., $M_{t+1}\leftarrow M_{\theta}$.
@@ 3 Experimental Setup
We evaluate the effectiveness of ScPO on a range of math and logical reasoning datasets:
@@
We compare models trained with ScPO in unsupervised (denoted as ScPOUnsup.) and semi-supervised (denoted as ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$) settings against the following:
@@
  • Language Models Self-Improved (LMSI). Following Huang et al. (2023), we implement LMSI, another unsupervised baseline that uses LLM self-consistency to generate target CoT solutions for problems and iteratively trains the LLM via supervised finetuning, i.e., the NLL loss, differing from ScPO’s weighted preference-based loss. Similar to ScPO, we generate additional reasoning problems using the LLM followed by consistency-based filtering (detailed in Section 2).
Table 1: GSM8K zero-shot accuracy after training Llama-3 Base 8B with ScPO and baselines, using greedy or 8-way self-consistency (SC)-based inference. The best performance is in bold, and second-best is underlined. We list train set sizes for each method: “Seed” corresponds to seed problems in the train set, whereas “Gen.” indicates additional problems generated by the model (without answers). IRPOGold, and ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$, highlighted in green, use the gold answers to create preference pairs (when available, indicated with ).
@@
llr@ / lcc Method Iter. Train Data (K) Test Acc. (%)
# Seed Gen. Greedy SC
@@
 IRPORM $M_{1}$ 5.5 - 48.67 69.98
$M_{2}$ 4.4 - 50.11 61.25
 LMSI $M_{1}$ 5.3 - 53.53 63.91
$M_{2}$ 1.1 5.2 56.71 62.55
 ScPOUnsup. $M_{1}$ 5.3 - 61.03 71.49
$M_{2}$ 1.4 5.1 63.91 71.11
@@
 IRPOGold $M_{1}$ 4.4 - 61.41 72.93
$M_{2}$ 5.7 - 64.29 72.56
 ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ $M_{1}$ 4.4 1.9 63.61 74.30
$M_{2}$ 5.7 4.5 66.64 74.75
@@ 4.1 Math Reasoning
Comparing methods on GSM8K, in Section 3, we observe that training with only one iteration of ScPO outperforms the zero-shot seed model and IRPORM, by $22.74\%$ and $12.36\%$, respectively, using greedy decoding. Similarly, on MATH (cf. Section 4.1), two iterations of ScPOUnsup. yields an improvement of $5.26\%$ and $1.64\%$ respectively compared to the same two baselines. We further note that while IRPORM is not given direct access to the gold labels, it uses the ArmoRM, which has been trained on human-annotated step-level data based on MATH’s train set (Lightman et al., 2024; Wang et al., 2024a). Hence, ScPO’s improvement over IRPORM would likely be larger if the RM had not used in-domain gold labels during training. Overall, we find ScPO has the ability to outperform RMs, especially in out-of-distribution settings. Lastly, in comparison to LMSI, another iterative and unsupervised baseline, two iterations of ScPOUnsup. outperform that of LMSI by $7.20\%$ and $2.76\%$ on GSM8K and MATH, respectively, when using greedy decoding. This highlights the importance of a weighted preference objective in training LLMs effectively using self-consistency.
Table 2: MATH zero-shot accuracy after training Llama-3 Base 8B with ScPO and baselines, using greedy or 8-way self-consistency (SC)-based inference. “Seed” corresponds to seed queries in the train set, “Gen.” are additional model-generated problems (without answers). IRPOGold and ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$, highlighted in green, use gold answers to train (indicated with ).
@@
llr@ / lcc Method Iter. Train Data (K) Test Acc. (%)
# Seed Gen. Greedy SC
@@
 IRPORM $M_{1}$ 6.4 - 18.06 24.20
$M_{2}$ 6.5 - 18.08 22.64
 LMSI $M_{1}$ 0.6 1.2 16.78 22.92
$M_{2}$ 1.1 2.0 16.96 20.20
 ScPOUnsup. $M_{1}$ 0.6 1.2 17.36 25.70
$M_{2}$ 1.2 2.5 19.72 24.58
@@
 IRPOGold $M_{1}$ 2.7 - 18.64 26.88
$M_{2}$ 3.0 - 20.32 26.88
 ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ $M_{1}$ 2.7 1.2 19.88 27.35
$M_{2}$ 3.0 2.2 20.48 26.92
Table 3: ZebraLogic test performance after unsupervised training of Llama-3 Instruct 8B with ScPO, compared to baselines. “Seed” corresponds to original puzzles in the test set, whereas “Gen.” indicates additional puzzles generated. Taken from the Leaderboard.
@@
$M_{1}$ w/ LMSI 0.4 / 2.0 16.2 51.1 2.6 45.8
$M_{2}$ w/ LMSI 0.4 / 2.0 16.8 53.6 2.5 46.9
$M_{1}$ w/ ScPOUnsup. 0.4 / 2.0 17.0 54.3 2.5 47.6
$M_{2}$ w/ ScPOUnsup. 0.5 / 2.2 18.1 58.2 2.5 45.2
Iterations of ScPO improve reasoning.
Unsupervised ScPO is comparable to IRPO training with gold labels.
We can compare the unsupervised training of ScPO with the supervised training using gold labels of IRPO in Sections 3 and 4.1. The results show that ScPOUnsup. without using any gold labels can yield comparable accuracy to IRPOGold on GSM8K and MATH with $<1\%$ gap in greedy performance and $<2\%$ gap in accuracy using 8-way self-consistency after two iterations of training ($M_{2}$). This comparable performance of ScPOUnsup. is likely due to high correlation (0.8 across the datasets) between the vote shares and accuracy on the test set, as further discussed in Appendix A. Note that on tasks that are challenging for the seed model $M_{0}$, such as MATH, we can only bootstrap a small set of examples from the original set of training problem as compared to IRPO (i.e., only around a quarter of examples obtain a clear majority answer). However, we can offset this gap in training data by generating new problems using few-shot prompting (cf. Section 2) and creating preference pairs using our self-consistency method. This yields improvements during the second iteration.
Semi-supervised training with ScPO outperforms IRPO. Lastly, in Sections 3 and 4.1, we evaluate the semi-supervised version of ScPO combined with using gold labels. We find that on GSM8K, ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ improves the greedy accuracy by $2.35\%$ and SC accuracy by $2.19\%$ in comparison to IRPOGold. Similar trends hold on the MATH dataset, where one iteration of ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ outperforms IRPOGold by $1.24\%$ using greedy decoding. These results show the utility of using ScPO to bootstrap from model-generated problems even with access to a labeled training set.
In Appendix C, we repeat the math reasoning experiments with Llama-3.1 Base 8B and find that while the absolute performance increases, the relative trends among the baselines remain the same – with two iterations of ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ improving the greedy test accuracy of the seed model by $25.32\%$ and $8.66\%$ on GSM8K and MATH, respectively.